# A Two-Stage Fourth-Order Implicit Scheme for Stiff problems

### v1

@@ 6 Conclusions and future work
This paper presents a novel two-stage fourth-order implicit scheme. We rigorously establish its fourth-order accuracy through mathematical analysis and systematically examine the scheme’s stability, deriving the conditions under which A-stability can be achieved. A detailed implementation procedure based on Newton iteration is provided. The effectiveness of the proposed method is validated through numerical experiments on several classical strongly stiff benchmark cases. Comparisons with classical explicit and implicit fourth-order Runge-Kutta methods demonstrate that the proposed scheme possesses superior stability regions and smaller numerical errors.
In future work, we will combine this implicit scheme with its explicit counterpart—a spatiotemporally coupled two-stage fourth-order method—to solve fluid dynamics problems characterized by strongly stiff source terms, such as those encountered in chemical reactions. Compared to other types of implicit schemes, the proposed method incorporates temporal derivatives of physical quantities, which is expected to enhance both computational accuracy and compactness of the scheme.
@@ 7 Acknowledgments
Zhixin Huo’s research work has been supported by the Key Program of Henan Higher Education Institutions (Grant No. 26A110007), the Young Talents Fund of Henan Province (Grant No. 252300423500), the "Double First-Class" Project of the School of Geomatics, Henan Polytechnic University (Grant No. BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No. B2024-60).

### v2

@@ 6 Conclusions and future work
This study fully leverages the time derivatives of physical quantities to construct a implicit TSFO temporal discretization scheme through rigorous mathematical derivation. A systematic theoretical analysis provides sufficient conditions for the scheme to achieve A-stability, and a corresponding Newton iteration process is proposed to accelerate convergence. The proposed algorithm is validated on a series of classical stiff benchmark problems, demonstrating its robustness and competitiveness. Numerical experiments show that the scheme achieves fourth-order temporal accuracy using only two stages. Compared with classical fourth-order implicit Runge-Kutta methods, it not only allows larger stable time steps but also yields smaller convergence errors.
This study establishes a solid theoretical foundation and a concrete implementation pathway for extending TSFO time discretization methods from existing explicit frameworks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] to implicit frameworks. Its significance and necessity lie not only in extending the applicability of TSFO methods from idealized homogeneous hydrodynamic problems to realistic ones with stiff source terms, but also in unlocking the inherent potential of Lax-Wendroff-based GRP solvers to naturally handle stiff source terms within the TSFO framework. The ability of TSFO methods to achieve fourth-order temporal accuracy with only two stages stems from their utilization of the time derivatives of physical quantities, which are precisely provided by Lax-Wendroff-based GRP solvers. Compared to traditional Riemann solvers, these advanced solvers can compute temporal derivatives that simultaneously integrate the physical effects of both fluxes and source terms within a unified framework through the governing equations, thereby effectively avoiding the numerical errors introduced by operator-splitting approaches.
Future work will integrate the existing explicit TSFO time discretization method [1] with the implicit TSFO scheme proposed in this paper, and further combine it with a more compact HWENO spatial discretization method based on GRP solver [5]. This integrated framework aims to solve hydrodynamic problems involving stiff source terms, targeting the simulation of complex, multi-scale, and multi-physics phenomena in real-world systems, such as chemically reacting flows, radiation hydrodynamics, combustion dynamics, and other stiff physical systems.
@@ 7 Acknowledgments
Zhixin Huo’s research work has been supported by the Key Program of Henan Higher Education Institutions (Grant No. 26A110007), the Young Talents Fund of Henan Province (Grant No. 252300423500), the Double First-Class Project of the School of Geomatics of Henan Polytechnic University (Grant No. BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No. B2024-60).